### v1

@@ 3.2 Parametric regularity for uniform random fields
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\!\!\!\!\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\!\!\!\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}$
$\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}.$
@@
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}$ $\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\cdot 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,|{\boldsymbol{\nu}}|\cdot\big(2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}\big)$
$\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\big(8+2E\,\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,e^{\lambda\,T}+16E\,\theta_{1}^{2}\,\theta_{2}\,e^{\lambda\,T}\big),$
where assumed $\rho\geq 1$ so that $\rho^{|{\boldsymbol{\nu}}|-3}\leq\rho^{|{\boldsymbol{\nu}}|-2}$, and used $D^{2}\leq DE$ (see (3.20)), together with
@@
$\rho:=2\,\Phi\,\big(4+E\,\theta_{1}^{2}\,\theta_{2}\,(\kappa_{\max}+8)\,e^{\lambda\,T}\big)\geq 1,$ (3.21)
@@
We have written the hypothesis (3.19) in a convenient form for induction, where the factor $\rho$ served as a scaling parameter that can be adjusted to suit our needs, as we defined in (3.21). Since $D=E\,\|w_{0}\|_{L^{2}(\Omega)}$ (see (3.20)), this factor $\rho$ can be seen as a rescaling of the sequence $\beta_{j}$ as follows:
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{D}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{E}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}}.$
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}w\|_{\mathcal{X}}\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|w_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}.$
@@ 4 QMC error analysis for uniform random fields
where the subtraction by ${\boldsymbol{\tfrac{1}{2}}}=(\tfrac{1}{2},\ldots,\tfrac{1}{2})$ translates points from the standard unit cube $0,1]^{s}$ to $U_{s}$. Alternatively, we may wish to compute the expectation of a QoI given by a bounded linear functional $G$ of the PDE solution $u_{s}^{\boldsymbol{y}}$, in which case we approximate ${\mathbb{E}}[G(u_{s})]\approx Q_{N}(G(u_{s}^{\boldsymbol{y}}))$, with the QMC rule applied analogously as in (4.1). For now we work with the exact PDE solution (i.e., no finite element method or time discretization).
@@
where $\widetilde{C}$ is from Theorem 3.1 and $C^{*}>0$ is independent of $s$.
@@ 5.1 Experiments with uniform random fields
The initial condition $u_{0}({\boldsymbol{x}})$ is a Gaussian function in the center of the domain as in [40] and the terminal time is taken to be $T=7$ days. Radiotherapy (2.3) is applied the first five days and chemotherapy (2.4) is administered daily. The radiotherapy dosage is uniformly $z_{\text{rt},k}({\boldsymbol{x}})=2$ Gy and the chemotherapy concentration is assumed to be uniformly $z_{\text{ct},\ell}({\boldsymbol{x}})=1$. The radiosensitivity parameter is $\alpha_{\text{rt}}=0.025\,\text{Gy}^{-1}$ with ratio $\alpha_{\text{rt}}/\beta_{\text{rt}}=10$ Gy following reported values in the literature [40]. The chemotherapy efficacy is $\alpha_{\text{ct}}=0.9$ with clearance rate $\beta_{\text{ct}}=24/1.8~\text{day}^{-1}$. The implicit Euler method is used to discretize (2.1) in time with time step $\Delta t=1/8$ day, such that $\gamma_{\text{rt},\varepsilon}=1/\Delta t$. Linear finite elements are used to discretize the spatial dimension with mesh size $h=1\,\text{mm}$.
@@
with stochastic dimension $s=256$, that is, there are $128$ terms in the expansion for each field. To obtain an unbiased estimator of ${\mathbb{E}}[G(u_{s}^{\boldsymbol{y}})]$, we average over $R$ i.i.d. uniform random shifts,
@@
where, similarly to (4.1), $Q^{(r)}_{N}[G(u_{s})]$ denotes the QMC approximation of the integral ${\mathbb{E}}[G(u_{s})]$ with the $r$th random shift. To estimate the root-mean-square error in Theorem 4.2, we compute the approximation
@@
Here we use $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[4,5,\dots,17]$, for a total of $2^{21}=2,097,152$ PDE evaluations. An off-the-shelf embedded lattice rule with weights tailored to a spectral decay rate of $1/j^{2}$ is utilized to generate the QMC sequence [11]. For comparison, we use $N^{\text{MC}}=2^{17}\cdot R=2^{21}$ quadrature points for MC to match the total number of PDE evaluations. Computations are carried out on the Texas Advanced Computing Center’s Frontera machine [41]. Code to reproduce the experiments is available at: https://github.com/gtpash/qmc-tumor. The results are displayed in Figure 1. We observe that the QMC root-mean-square error converges at a linear rate ${\mathcal{O}}(N^{-1})$ while standard MC converges with the typical $\mathcal{O}(N^{-1/2})$ rate, consistent with the theory.

### v2

@@ 3.2 Parametric regularity for uniform random fields
$\quad+\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+\theta_{1}^{2}\,\theta_{2}\sum_{j\geq 1}\nu_{j}\,\beta_{j}\!\!\!\!\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\!\!\!\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}D\,{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\rho^{|{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}}\rho^{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}$
$\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}<{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,\sum_{j\geq 1}\nu_{j}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}\binom{{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{m}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{m}}|}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}-{\boldsymbol{e}}_{j}-{\boldsymbol{m}}|}.$
@@
$(1/\Phi)\,\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}$ $\leq 2\,D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,|{\boldsymbol{\nu}}|\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+D^{2}\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\cdot 2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}$
$\quad+D^{2}\theta_{1}^{2}\,\theta_{2}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-3}\,|{\boldsymbol{\nu}}|\cdot\big(2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}+2\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|-1}\big)$
$\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-2}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\big(8+2E\,\theta_{1}^{2}\,\theta_{2}\,\kappa_{\max}\,|\Omega|^{1/2}+16E\,\theta_{1}^{2}\,\theta_{2}\,|\Omega|^{1/2}\big),$
where we assumed $\rho\geq 1$ so that $\rho^{|{\boldsymbol{\nu}}|-3}\leq\rho^{|{\boldsymbol{\nu}}|-2}$, and used $D^{2}\leq DE\,|\Omega|^{1/2}$ (see (3.22)), together with
@@
$\rho:=2\,\Phi\,\big(4+E\,\theta_{1}^{2}\,\theta_{2}\,(\kappa_{\max}+8)\,|\Omega|^{1/2}\big)\geq 1,$ (3.23)
@@
We have written the hypothesis (3.21) in a convenient form for induction, where the factor $\rho$ served as a scaling parameter that can be adjusted to suit our needs, as we defined in (3.23). Since $D=E\,\|u_{0}\|_{L^{2}(\Omega)}$ (see (3.22)), this factor $\rho$ can be seen as a rescaling of the sequence $\beta_{j}$ as follows:
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq D\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\rho^{|{\boldsymbol{\nu}}|-1}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{D}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}=\tfrac{E}{\rho}\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}\neq{\boldsymbol{0}}.$
@@
$\|\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}u\|_{\mathcal{X}}\leq\max(C,\tfrac{E}{\rho})\,(\rho\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}\,{[\tfrac{1}{2}]}_{|{\boldsymbol{\nu}}|}\,\|u_{0}\|_{L^{2}(\Omega)}\quad\mbox{for all }{\boldsymbol{\nu}}.$
@@ 4 QMC error analysis for uniform random fields
As noted in the introduction, QMC methods have demonstrated impressive performance in several UQ applications involving PDEs with random coefficients or domains, including linear elliptic PDEs [13, 17, 20, 21, 22, 24, 32], elliptic eigenvalue problems [19, 40] and linear parabolic equations with control [23]. For a more pedagogical survey of some of the earlier works from the QMC perspective, we refer the interested reader to [31].
@@
where the subtraction by ${\boldsymbol{\tfrac{1}{2}}}=(\tfrac{1}{2},\ldots,\tfrac{1}{2})$ translates points from the standard unit cube $0,1]^{s}$ to $U_{s}$. Alternatively, we may wish to compute the expectation of a QoI given by a bounded linear functional $G$ of the PDE solution $u_{s}^{\boldsymbol{y}}$, in which case we approximate ${\mathbb{E}}[G(u_{s})]\approx Q_{N}(G(u_{s}^{\boldsymbol{y}}))$, with the QMC rule applied analogously as in (4.1). A randomly shifted lattice rule is unbiased, i.e., ${\mathbb{E}}_{\boldsymbol{\Delta}}[Q_{N}(G(u_{s}))]={\mathbb{E}}[G(u_{s})]$. For now we work with the exact PDE solution (i.e., no finite element method or time discretization).
@@
where $\widetilde{C}$ is the constant from Theorem 3.1 and $C^{*}>0$ is independent of $s$.
@@ 5.1 Experiments with uniform random fields
The initial condition $u_{0}({\boldsymbol{x}})$ is a Gaussian function in the center of the domain as in [44] and the terminal time is taken to be $T=7$ days. Radiotherapy (2.3) is applied the first five days and chemotherapy (2.4) is administered daily. The radiotherapy dosage is uniformly $z_{\text{rt},k}({\boldsymbol{x}})=2$ Gy and the chemotherapy concentration is assumed to be uniformly $z_{\text{ct},\ell}({\boldsymbol{x}})=1$. The radiosensitivity parameter is $\alpha_{\text{rt}}=0.025\,\text{Gy}^{-1}$ with ratio $\alpha_{\text{rt}}/\beta_{\text{rt}}=10$ Gy following reported values in the literature [44]. The chemotherapy efficacy is taken to be $\alpha_{\text{ct}}=0.9$ with clearance rate $\beta_{\text{ct}}=24\,\ln(2)/1.8~\text{day}^{-1}$ for a chemotherapy based on Temozolomide, which has a half-life of 1.8 hours [16]. The implicit Euler method is used to discretize (2.1) in time with time step $\Delta t=1/8$ day, such that $\gamma_{\text{rt},\varepsilon}=1/\Delta t$, where $\varepsilon=\Delta t$ matches the resolution of the time-discretization. From (2.3), for each $k$, we compute the time-integrated effect of one radiotherapy dose
$\int f_{\text{rt},k}\,dt=1-\exp(-\alpha_{\text{rt}}z_{\text{rt},k}-\beta_{\text{rt}}z^{2}_{\text{rt},k})\approx 0.06,$
hence each dose of radiotherapy reduced $u$ by about $6\%$. Similarly, for each $\ell$,
$\int f_{\text{ct},\ell}\,dt=\alpha_{\text{ct}}z_{\text{ct},\ell}\int_{\tau_{\text{ct},\ell}}^{\infty}\exp(-\beta_{\text{ct}}(t-\tau_{\text{ct},\ell}))=\frac{\alpha_{\text{ct}}\,z_{\text{ct},\ell}}{\beta_{\text{ct}}}\approx 0.10,$
and so each dose of chemotherapy reduces $u$ by about $10\%$, with a slight time delay. Linear finite elements are used to discretize the spatial dimension with mesh size $h=1\,\text{mm}$.
@@
with stochastic dimension $s=256$, that is, there are $128$ terms in the expansion for each field. To obtain an unbiased estimator of ${\mathbb{E}}[G(u_{s})]=\int_{U_{s}}G(u_{s}^{\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}$, we average over $R$ i.i.d. uniform random shifts,
@@
where, similarly to (4.1), $Q^{(r)}_{N}[G(u_{s})]$ denotes the unbiased QMC approximation of the integral ${\mathbb{E}}[G(u_{s})]$ with the $r$th random shift. Therefore the average $\overline{Q}$ is also an unbiased estimator of ${\mathbb{E}}[G(u_{s})]$, and its variance is $1/R$ times the variance for any individual $Q^{(r)}_{N}[G(u_{s})]$. To estimate the root-mean-square error in Theorem 4.2, we compute the approximation
@@
Here we use $R=16$ random shifts $\boldsymbol{\Delta}$} and $N=2^{m}$ lattice points with $m=[4,5,\dots,17]$, for a total of $2^{21}=2,097,152$ PDE evaluations. An off-the-shelf embedded lattice rule with weights tailored to a spectral decay rate of $1/j^{2}$ is utilized to generate the QMC sequence [11]. For comparison, we use $N^{\text{MC}}=2^{17}\cdot R=2^{21}$ quadrature points for MC to match the total number of PDE evaluations. Computations are carried out on the Texas Advanced Computing Center’s Lonestar6 cluster across two compute nodes with 256 total parallel MPI processes, each evaluating one PDE. The MC reference calculation with $2^{21}$ evaluations took 1,141 minutes ($\approx 19$ hours), while the QMC shift replicates with $2^{17}$ evaluations averaged 146 minutes ($\approx 2.4$ hours). Code to reproduce the experiments is available at: https://github.com/gtpash/qmc-tumor. The results are displayed in Figure 1. We observe that the QMC root-mean-square error converges at a linear rate ${\mathcal{O}}(N^{-1})$ while standard MC converges with the typical $\mathcal{O}(N^{-1/2})$ rate, consistent with the theory.